OpenStudy (anonymous):

If Y is the midpoint of TW, TY = 4x - 4 and YW = 2x + 8, find TY. A. 5 B. 6 C. 14 D. 20

OpenStudy (anonymous):

i think the problem is asking about the intersection point of the graphs of these functions

OpenStudy (anonymous):

Since Y is a midpoint, TY=YW. Therefore, 4x-4=2x+8. Solving this linear equation we get: 2x=12 x=6. Since TY=4x-4, TY=4*6-4=20